an airplane made a flight of 1600 miles in 5 hours. during the first 3 hours of the trip it had good weather. it then ran into bad weather, which decreased its rate by 75 mph for the rest of the flight. find the rate of each part of the flight.
Answer by Mathtut(3670)

lets call
x the speed of the aircraft the first 3 hours and
x-75 the speed of the aircraft the final 2 hours.
so: 3x + 2(x-75) = 1600

distribute the 2nd term

: 3x + 2x-150=1600 combine like terms
: 5x=1750 x=350
so the rate of first 3 hours was 350mph
and the rate of the final 2 is 350-75=275mph